OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4)

OpenStudy (nubeer):

Try using this formula.. \[y-y _{1}= m(x- x _{1})\]

OpenStudy (anonymous):

Use this formula y-y1 =m(x-x1) \[y +4 = \frac{ -1 }{ 3 }(x-6)\] sovle for y.

OpenStudy (nubeer):

m = slope. x1 and y1 are coordinates..

OpenStudy (anonymous):

Yeah I know to use the formula I just wanted to confirm I was right

OpenStudy (anonymous):

Your right.